# HOI-Swap: Swapping Objects in Videos with Hand-Object Interaction Awareness

### v1

@@ B.2 Baselines
We implement the baselines described in Section 4.1 using their official code. AnyDoor is recognized as the best image editing method and is integrated into the two video editing baselines, per-frame and AnyV2V. Note that these methods impose more demanding test-time requirements compared to ours. PBE and AnyDoor both necessitate precise segmentation masks of the source object, while HOI-Swap only requires a bounding box during inference. In addition, VideoSwap requires users to click semantic points for every video, provide a text prompt describing the video, and train a specialized object model using a few reference images. This extensive prepossessing limits its scalability and is a contributing factor to the smaller size of our video evaluation set. In contrast, HOI-Swap is fast on inference, with RAFT optical flow extraction completing in just a few seconds per video, making it efficient and practical for real-world applications.
@@ B.4 Implementation
For data preprocessing, we apply several transformations to the reference object image: random rotations with angles ranging from -90 to 90 degrees, random perspective changes, and random horizontal flips. We crop and center the source frame using a variable ratio between 0.3 and 0.6, determined during training to introduce variability; this ratio is defined as the size of the bounding box relative to the image size. For stage-I training: image resolution is set as 512$\times$512. We train the model for a total of 25K steps with a learning rate of 1e-4 and a batch size of 8. We finetune the entire 2D UNet for image editing. For stage-II training: input video resolution is set as 14$\times$256$\times$256, where we sample 14 frames at an fps of 7. The model is trained for a total of 50K steps with a learning rate of 1e-5 and a batch size of 1. We finetune the temporal layers of the 3D UNet. A classifier-free guidance dropout rate of 0.2 is employed for all stages. Training for each stage takes about 3 days on one 8-NVIDIA-V100-32G GPU node.
@@ C.1 More qualitative results
Beyond the examples shown in Figure 4 of the main paper, we provide additional qualitative image editing results of HOI-Swap in Figures 8 and 9. HOI-Swap excels in both structured scenarios with relatively clean backgrounds (Figure 8, source frames from HOI4D and TCN Pouring), and challenging in-the-wild scenarios featuring cluttered backgrounds (Figure 9, source frames from EgoExo4D and EPIC Kitchens). Compared to the baselines, HOI-Swap seamlessly integrates the reference object into the target scene, demonstrating exceptional HOI awareness and the ability to accurately reorient the reference object to align with the scene context. Moreover, it also performs well in general object-swapping settings where no hand is involved, as shown in rows 3, 5, and 6 of Figure 9. Please refer to Supp. video for more qualitative results of video editing.
@@
C.2 User study analysis
@@ C.3 Ablation study
Figure 11: Ablation study of sampled motion points sparsity. The left figure illustrates a scenario where only partial motion transfer is desired, due to differences between the original and new object. The right figure showcases a scenario where full motion transfer is beneficial, owing to the similarities between the objects. We invite readers to view these examples in Supp. video.
@@
C.4 Discussion on sampling region
@@
C.5 Comparison with text-guided editing approaches
@@ Appendix D Limitations
In addition, we present failure cases in the Supp. video, stemming from stage-I failure (where the model does not establish accurate spatial alignment) and stage-II failure (where the model fails to correctly understand the reference object as a whole). We invite readers to review these examples for a better understanding of the potential issues.

### v2

@@ 6 Conclusion
Acknowledgements: UT Austin is supported in part by the IFML NSF AI Institute. KG is paid as a research scientist at Meta. The authors would like to thank Zhengqi Gao (MIT), Xixi Hu (UT Austin), Hanwen Jiang (UT Austin), Feng Liang (UT Austin), and Yue Zhao (UT Austin) for their insightful discussions and valuable feedback, which contributed to the development of this work.
@@ B.2 Baselines
We implement the baselines described in Section 4.1 using their official code. Baselines including PBE, AnyDoor and AnyV2V are adopted in a zero-shot manner. Afford Diff utilizes the same training data source (HOI4D) as ours. We do not provide additional layout or hand orientation information to Afford Diff. VideoSwap [17], a one-shot-tuned video editing approach, is trained on the source video using the default setup in their official repository. AnyDoor is recognized as the best image editing method and is integrated into the two video editing baselines, per-frame and AnyV2V.
Note that these baseline methods impose more demanding test-time requirements compared to ours. PBE and AnyDoor both necessitate precise segmentation masks of the source object, while HOI-Swap only requires a bounding box during inference. In addition, VideoSwap requires users to click semantic points for every video, provide a text prompt describing the video, and train a specialized object model using a few reference images. This extensive prepossessing limits its scalability and is a contributing factor to the smaller size of our video evaluation set. In contrast, HOI-Swap is fast on inference, with RAFT optical flow extraction completing in just a few seconds per video, making it efficient and practical for real-world applications.
@@ B.4 Implementation
For data preprocessing, we apply several transformations to the reference object image: random rotations with angles ranging from -90 to 90 degrees, random perspective changes, and random horizontal flips. We crop and center the source frame using a variable ratio between 0.3 and 0.6, determined during training to introduce variability; this ratio is defined as the size of the bounding box relative to the image size. For stage-I training: image resolution is set as 512$\times$512. We train the model for a total of 25K steps with a learning rate of 1e-4 and a batch size of 8. We finetune the entire 2D UNet for image editing. For stage-II training: input video resolution is set as 14$\times$256$\times$256, where we sample 14 frames at an fps of 7. The model is trained for a total of 50K steps with a learning rate of 1e-5 and a batch size of 1. We finetune the temporal layers of the 3D UNet. A classifier-free guidance dropout rate of 0.2 is employed for all stages. Training for each stage takes about 3 days on one 8-NVIDIA-V100-32G GPU node. We set sampling points sparsity value as 50% for all quantitative evaluations.
@@ C.1 More qualitative results
Beyond the examples shown in Figure 4 of the main paper, we provide additional qualitative image editing results of HOI-Swap in Figures 9 and 10. HOI-Swap excels in both structured scenarios with relatively clean backgrounds (Figure 9, source frames from HOI4D and TCN Pouring), and challenging in-the-wild scenarios featuring cluttered backgrounds (Figure 10, source frames from EgoExo4D and EPIC Kitchens). Compared to the baselines, HOI-Swap seamlessly integrates the reference object into the target scene, demonstrating exceptional HOI awareness and the ability to accurately reorient the reference object to align with the scene context. Moreover, it also performs well in general object-swapping settings where no hand is involved, as shown in rows 3, 5, and 6 of Figure 10. Please refer to our project page for more qualitative results of video editing.
@@
C.2 Detailed quantitative results
Table 1 of the main paper reports video editing results when videos are split based on object instances. In addition, we conduct experiments with two splitting ways: (1) across different subjects, and (2) across different actions. The results are reported in Table 2. HOI-Swap consistently outperforms baseline approaches across different data splitting settings, demonstrating its effectiveness.
Furthermore, we provide a breakdown of video editing results in Table 3, separating results from in-domain videos (HOI4D) and out-of-domain videos (EPIC-Kitchens & TCN Pouring). We observe that the performance gain of HOI-Swap is higher for in-domain videos than out-of-domain videos. Future work could explore extending HOI-Swap’s generalization capabilities.
Table 2: Video editing results with splitting by subject and action (Table 1 in the main paper reports results with splitting by object instances).
Method split by subject split by action
subj. cons. mot. smth. cont. agr. hand agr. hand fid. subj. cons. mot. smth. cont. agr. hand agr. hand fid.
Per-frame [8] 85.6 95.2 73.6 63.3 62.9 84.5 94.7 82.8 57.6 71.0
AnyV2V [29] 67.2 95.3 68.2 27.7 10.0 69.8 94.7 74.6 30.6 19.9
VideoSwap [18] 88.0 97.4 71.8 29.6 69.0 87.2 97.1 87.7 39.8 79.4
HOI-Swap (ours) 90.0 98.2 80.4 84.4 90.3 89.6 97.8 93.3 77.4 94.6
Table 3: Video editing results breakdown: in-domain videos (left) and out-of-domain videos (right).
Method In-domain videos Out-of-domain videos
subj. cons. mot. smth. cont. agr. hand agr. hand fid. user pref. subj. cons. mot. smth. cont. agr. hand agr. hand fid. user pref.
Per-frame [8] 85.8 94.7 81.3 61.5 78.4 0.3 91.0 96.4 97.3 71.9 90.2 0.0
AnyV2V [29] 73.1 95.5 70.9 33.2 27.0 1.7 68.8 95.7 62.7 25.6 26.9 0.0
VideoSwap [18] 90.5 97.5 82.3 38.9 74.8 0.4 94.4 98.1 99.6 81.5 97.1 2.5
HOI-Swap (ours) 91.3 98.0 89.9 79.0 96.6 82.7 94.7 98.6 99.8 77.8 99.5 70.2
C.3 User study analysis
@@ C.4 Ablation study
DINO vs. CLIP encoder
Stage I and stage II of our pipeline employ the DINO and CLIP encoders, respectively, to align with their specific objectives. Stage I, focused on swapping the reference object within a single frame, benefits from the DINO encoder due to its enhanced ability to capture “objectness” compared with the CLIP encoder. The main emphasis of stage II is to transfer motion from the source video, and a CLIP encoder is adopted to provide scene context for generating the video. We additionally conduct experiments to explore the possibility of using a DINO encoder for stage II. As reported in Table 5, the two encoder variants perform similarly.
Table 5: Video editing results: comparison of DINO and CLIP encoders for stage II. The two object encoders yield similar performance.
Video Editing subject consistency motion smoothness contact agreement hand agreement hand fidelity
HOI-Swap (DINO encoder) 91.2 98.1 88.9 79.9 96.2
HOI-Swap (CLIP encoder) 91.4 98.0 89.9 79.0 96.6
Ground truth vs. estimated object masks
While it is generally assumed that users will provide the ground truth bounding box or segmentation mask of the object they wish to replace during inference [57, 7], we acknowledge the potential of incorporating automatic segmentation methods to reduce user effort. Following this, we applied SAM-2 [44] to identify bounding boxes on test videos as an alternative to manually providing ground truth. This method requires just a single click inside the object in the initial frame, followed by SAM-2 automatically tracking the target object. The quantitative comparisons are presented in Table 6. While there is some degradation across three HOI metrics, we believe this feature is valuable for downstream applications as it greatly eases the input requirements and improves user convenience. Note that the baseline approaches have more stringent input requirements than HOI-Swap, e.g. need precise object segmentation masks or additional text prompts. Even with the use of automatically generated masks, HOI-Swap demonstrates its great advantages over the baselines.
Table 6: Video editing results: comparison of using ground truth object bounding boxes vs. SAM-2 estimated ones as model input.
Video Editing subject consistency motion smoothness contact agreement hand agreement hand fidelity
Prior best 90.5 97.5 82.4 61.5 78.4
HOI-Swap (SAM-2 bbox) 91.2 98.0 87.8 73.9 91.8
HOI-Swap (GT bbox) 91.4 98.0 89.9 79.0 96.6
Masking strategy
In stage I, we mask the source frame with a square bounding box, as opposed to using the source object’s segmentation mask as in  [57]. This masking strategy not only directs the model to fill the predefined masked area but also to generate plausible hand-object interactions aligning with the source frame’s hand position and the reference object. To demonstrate this point, we experimented with a variant of the stage I model that uses the original object’s segmentation mask instead. As illustrated in Figure  12, this variant struggles with grasp changes when the reference and source objects differ, thereby reinforcing the effectiveness of our chosen approach.
Figure 12: Comparison of two masking strategies. The stage-I model using a segmentation mask (row 2) only learns to fill the missing region without HOI awareness. In contrast, a square bounding box (row 3) also masks the hand-object interaction regions, requiring the model to reconstruct these interactions; this enables our stage-I model to effectively learn diverse HOI patterns from data.
@@
Figure 13: Ablation study of sampled motion points sparsity. The left figure illustrates a scenario where only partial motion transfer is desired, due to differences between the original and new object. The right figure showcases a scenario where full motion transfer is beneficial, owing to the similarities between the objects. We invite readers to view these examples in our project page.
@@
C.5 Discussion on sampling region
@@
C.6 Comparison with text-guided editing approaches
@@ Appendix D Limitations
Figure 16: Failure cases of HOI-Swap. Left: stage-I failure, Right: stage-II failure.
@@
In addition, we present failure cases in Figure 16. The left example shows a stage-I failure, where HOI-Swap fails to achieve accurate spatial alignment. The right example illustrates a stage-II failure, where noise in the warped sequence impedes model’s understanding of the reference object, leading to wrong motion patterns.